What is the [H3O+] of a solution with pOH= (5.30x10^0)? Enter your answer in scientific notation and three sig figs.
Answer #1

We know pH + pOH = 14.

Given pOH = 5.3*10^0 = 5.3 since 10^0=1.

So, pH = 14-pOH = 14-5.3 = 8.7

Now, pH is defined as the negative log of Hydronium ion concentrations.

So, pH = - log( H3O+)

8.7 = -log (H3O+), or H3O+ = 10^-8.7

= ​​​​​​​​​ ​​​​​1. 995*10-9 M (Answer) concentration of Hydronium ion